OpenStudy (anonymous):

Hello guys, I would love it if you could help me! I am currently in 6th grade and there's no one else who can help me with my homework besides well.... You guys! Ok so my 3 maths questions I need help with are -3/9 = -1/2 2x/3=4/5 4a+1=3 Please tell me how to work it out! I'm learning about basic equations! Thanks in general

OpenStudy (wolf1728):

Are those numbers two equations? Also I'd like to know exactly how those equations are stated? Is there an online source you can copy and paste?

OpenStudy (anonymous):

Do you have skype or something that I could send a picture? sorry for bothering you!

OpenStudy (wolf1728):

I don't have Skype You can attach a picture or text to this topic.

OpenStudy (anonymous):

How so, I am on a mobile device! Maybe Facebook? Do you have any sort of social media, if not, it's an equation

OpenStudy (wolf1728):

I don't want you to go to too much trouble. Anyway, the equations are: -3/9 = -1/2 2x/3=4/5 4a+1=3 Do I have that right?

OpenStudy (anonymous):

Yes! Thank you!

OpenStudy (wolf1728):

I'm having a LOT of trouble with Open Study at the moment - it keeps kicking me off

OpenStudy (anonymous):

My mistake I meant for the fit's question is 3/a

OpenStudy (anonymous):

first*

OpenStudy (wolf1728):

so it is -3/a = -1/2

OpenStudy (anonymous):

Sure Is!!

OpenStudy (wolf1728):

Do you know how to begin solving for a ?

OpenStudy (anonymous):

At class, I was not taught this as we had a test, yet our teacher assigned us with this.

OpenStudy (anonymous):

So I'm kinda..... Clueless...

OpenStudy (wolf1728):

-3/a = -1/2 We can multiply both sides of the equation by "a" which gives us -3a/a = -1a/2 which can be neatened into: -3 = -a/2 Then we can multiply both sides by 2 giving us -3*2 = -a Do you see that?

OpenStudy (anonymous):

What does the star represent?

OpenStudy (wolf1728):

Multiply (it means -3 times 2)

OpenStudy (anonymous):

So for 3a/a, you cancel the 'a's?

OpenStudy (wolf1728):

sure it is just like saying 3*2 / 2 equals 3

OpenStudy (anonymous):

Ohhhh I get it!!! And answer even matches! Thanks! What about 2x/3 = 4/5?? Is it the exact same method?

OpenStudy (wolf1728):

Well we still have to finish solving for "a" -3 * 2 = -a

OpenStudy (anonymous):

so -3 * 2 = -6

OpenStudy (wolf1728):

Yes, -6 = -a so "a" equals?

OpenStudy (anonymous):

6???

OpenStudy (anonymous):

Because 2 minuses together equal a positive right?

OpenStudy (wolf1728):

Actually, the proper way to say it we multiply both sides by -1 and the equation becomes 6 = a but yes you got it right !

OpenStudy (anonymous):

Thanks!!! I think I've got this question now!

OpenStudy (anonymous):

So are we getting onto the next question?

OpenStudy (anonymous):

Or not

OpenStudy (wolf1728):

Okay the other 2 equations are 2x/3=4/5 4a+1=3 Let's try 2x/3=4/5 We can multiply both sides by 3 to get 3*(2x/3) = 3*(4/5) 2x = 12/5 Then we divide both sides by 2 x = 6/5

OpenStudy (anonymous):

Oh sorry hah! I can't see whether you are typing or not! Weird glitch

OpenStudy (wolf1728):

Yeah Open Study is giving me a LOT of trouble at the moment. Anyway moving on to 4a+1=3 we subtract 1 from both sides 4a = 3 -1 4a = 2 Next we divide both sides by 4 and can you figure what "a" is?

OpenStudy (anonymous):

2??

OpenStudy (wolf1728):

No 4a/4 = 2/4 Does that help?

OpenStudy (anonymous):

Ahh yeah it does now!i get it and I've awarded you with a medal for you putting up with how dumb I am. I know, I shouldn't even be asking here, this is for advanced maths so I should probably delete this post after I write down the formulas!

OpenStudy (wolf1728):

Heck there is always somebody to help anyone else no matter if it is beginning algebra or advanced calculus. Thatks for the medal and I'm glad I could help you out. I'd better leave Open Study - it is giving me a LOT of trouble.